# Improving Multimodal Learning Balance and Sufficiency through Data Remixing

### v1

@@ 4.5.4 Study of Unimodal Prediction Methods
Since our method requires accurate unimodal predictions, we add a classification head to each modality branch and synchronize the updates of the classification head parameters by modifying the loss function. This approach has been shown to improve model performance (wang2020makes). To further validate our method, we compare the results with the dropout method (dropout; wei2024enhancing) for unimodal prediction evaluation across three traditional fusion strategies. The results are presented in Table 6. From the table, we observe that both methods of obtaining unimodal predictions show effective improvements when applying our strategy. Notably, regardless of the chosen method, our approach does not introduce any additional overhead during inference.
@@ 5 Conclusion
Multimodal models often face challenges due to the differences in optimization trajectories between modalities, leading to issues of insufficient and imbalanced learning during joint training. We propose the Data Remixing method, which decouples multimodal data based on unimodal separability and reassembles unimodal data to ensure consistency between modalities at the batch level. Our method effectively mitigates both modality laziness and modality clash, achieving significant performance improvements across various datasets, fusion methods, and model structures. Additionally, it does not require dataset expansion or introduce extra computational overhead during inference. Moreover, to the best of our knowledge, we are the first to propose that the challenges of multimodal learning originate at the batch level and to offer a solution for them. However, a limitation of our approach arises when one modality serves primarily as an auxiliary modality with limited information (wei2025diagnosing). In such cases, the unimodal evaluation and allocation strategy may require further refinement.

### v2

@@ 4.5.4 Study of Unimodal Prediction Methods
Since our method requires accurate unimodal predictions, we add a classification head to each modality branch and synchronize the updates of the classification head parameters by modifying the loss function. This approach has been shown to improve model performance (Wang et al., 2020). To further validate our method, we compare the results with the dropout method (Hinton, 2012; Wei et al., 2024) for unimodal prediction evaluation across three traditional fusion strategies. The results are presented in Table 6. From the table, we observe that both methods of obtaining unimodal predictions show effective improvements when applying our strategy. Notably, regardless of the chosen method, our approach does not introduce any additional overhead during inference.
@@ 5 Conclusion
Multimodal models often face challenges due to the differences in optimization trajectories between modalities, leading to issues of insufficient and imbalanced learning during joint training. We propose the Data Remixing method, which decouples multimodal data based on unimodal separability and reassembles unimodal data to ensure consistency between modalities at the batch level. Our method effectively mitigates both modality laziness and modality clash, achieving significant performance improvements across various datasets, fusion methods, and model structures. Additionally, it does not require dataset expansion or introduce extra computational overhead during inference. Moreover, to the best of our knowledge, we are the first to propose that the challenges of multimodal learning originate at the batch level and to offer a solution for them. However, a limitation of our approach arises when one modality serves primarily as an auxiliary modality with limited information (Wei et al., 2025). In such cases, the unimodal evaluation and allocation strategy may require further refinement.